# Jet schemes, Newton polygons and continued fractions

### v2

@@ 8. Generating Series
In this section, we introduce the generating series associated to $f\in K[[x,y]]$ a Newton non-degenerate power series. Denote by $\mathcal{C}=\{f=0\}$ its associated plane curve singularity. This series will be denoted $G:=G(f)$.
@@
where $m\geq 1$ and $K$ is an irreducible component in $\text{Cont}^{m}(\mathcal{C})^{0}\neq\varnothing$. We call a subsummation of a sum $S$ to be a sum whose terms form a subset of the terms of $S$.
@@
  • $H$ is the subsummation of $G$ with $K$ being of the form $\mathcal{H}^{\beta}_{\infty}$;
  • $R$ is the rest of the sum since $G=H+R$.
We used the letter $H$ because the $K$s appearing in $H$ are all irreducible components having hyperplane coordinates. R=GHR=G-H is just the rest of the sum.
Let’s consider two different examples at first.
@@
Let’s start with the example of $f=(y^{2}-x^{3})(y^{2}-2x^{3})$. Its tropical variety $T(f)=\mathbf{R}_{\geq 0}(2,3)$ is a semi-line. Set $G=G(f)$. Introduce
@@
where $m\geq\nu_{\alpha}(f)$ and $\alpha\in T(f)$. Then, since $\alpha=s\cdot(2,3)$ for $s\in\mathbf{N}$,
$R=2u\sum_{s=1}^{\infty}\sum_{m=12s}^{\infty}u^{5s+m-12s}v^{m}=\frac{2u}{1-uv}\sum_{s=1}^{\infty}u^{5s}v^{12s}=\frac{2u}{1-uv}\frac{u^{5}v^{12}}{1-u^{5}v^{12}}$
Now, set $S_{1}=(0,4)$ and $S_{2}=(6,0)$. Introduce $S^{*}_{i}$ the corresponding cone to $S_{i}$ in the dual Newton fan. Set
@@
where $\beta_{i}\in B^{*}_{m}\cap\overset{\circ}{S^{*}_{i}},m\geq 1$. Let’s describe where $\beta_{i}$ lies.
@@
for $\lambda\in T(f)\cap\mathbf{N}_{0}^{2}$. So we get, $\beta_{2}-\lambda\in\{(1,2)\}$ for the same $\lambda=(2s,3s)$.
@@
Then, $H\in\mathbf{Z}(u,v)$. So, $G=H+R\in\mathbf{Z}(u,v)$.
@@
Consider $f=f_{1}f_{2}=(y^{2}-x^{3})(y^{3}-x^{5})$. Its tropical variety is
$T(f)=\mathbf{R}_{\geq 0}(2,3)\cup\mathbf{R}_{\geq 0}(3,5)$
Let’s compute
@@
where $\alpha_{i}\in T(f_{i})$, $m_{i}\geq\nu_{\alpha_{i}}(f)$.
$R_{1}=u\sum_{s=1}^{\infty}\sum_{m=0}^{\infty}u^{5s+m}v^{m+15s}=\frac{u}{1-uv}\frac{u^{5}v^{15}}{1-u^{5}v^{15}}$
$R_{2}=\frac{u}{1-uv}\frac{u^{8}v^{24}}{1-u^{8}v^{24}}$
Denote now, $S_{1}=(0,5)$, $S_{2}=(3,3)$ and $S_{3}=(8,0)$ the vertices of the Newton polygon from left to right. Introduce again $S_{i}^{*}$ the corresponding cone to $S_{i}$ in the dual Newton fan. Let’s compute
@@
where $\beta\in B^{*}_{m},m\geq 1$ and $\sigma_{2}=\mathbf{R}_{\geq 0}(2,3)+\mathbf{R}_{\geq 0}(3,5)=S^{*}_{2}$. We obtain by computing geometric series
$D:=H-\tilde{H}=\frac{u^{5}v^{15}}{1-u^{5}v^{15}}\frac{u^{8}v^{24}}{1-u^{8}v^{24}}\in\mathbf{Z}(u,v)$
Now, let’s compute
@@
where $H^{*}$ is the subsummation of $H$ whose $\beta$ lie in the first and last cone determined by orientation of the tropical decomposition.
Then, $\tilde{H}-H^{*}=\frac{u^{5}v^{15}}{1-u^{5}v^{15}}+\frac{u^{8}v^{24}}{1-u^{8}v^{24}}\in\mathbf{Z}(u,v)$. So, we have
@@
where $\beta_{i}\in\overset{\circ}{S_{i}^{*}}\cap B^{*}_{m},m\geq 1$, $S_{1}=(0,5)$ and $S_{3}=(8,0)$. Then,
@@
This leads to
$H^{*}=\frac{u^{2}v^{5}+u^{4}v^{10}}{1-u^{5}v^{15}}+\frac{u^{3}v^{8}+u^{6}v^{16}}{1-u^{8}v^{24}}\in\mathbf{Z}(u,v)$
So, $H\in\mathbf{Z}(u,v)$ and $G=H+R$ as well.
In the below lemma, we write an expression for $R$ (which was introduced at the beginning of the section).
@@
Let $f\in K[x,y]$ be Newton non-degenerate whose tropical decomposition is written as $f=f_{1}...f_{t}$. For $1\leq i\leq t$, write $f_{i}=\prod_{j=1}^{r_{i}}f_{i,j}$. Then
@@
where $1\leq i\leq t$, $1\leq j\leq r_{i}$, $\alpha\in T(f_{i})$ and $m\geq\nu_{\alpha}(f)$. Also, $R\in\mathbf{Z}(u,v)$ and
@@
Set $R_{i}=\sum_{j,\alpha_{i},m}u^{1+|\alpha_{i}|+m-\nu_{\alpha_{i}}(f)}v^{m}$ in such a way that $R=\sum_{i}R_{i}$. Since $\nu_{\alpha}(f)=\alpha\cdot S_{i}$ (because $\alpha\in T(f_{i})$). Thus, we obtain:
@@
Now let $f=f_{1}...f_{t}$ be a tropical decomposition of a Newton non-degenerate polynomial. The decomposition $f=f_{1}...f_{t}$ is oriented when for all $i<j$,
@@
ie $T(f_{j})$ is ”above” $T(f_{i})$ which we write $T(f_{j})>T(f_{i})$. Now let’s apply this new notion of oriented tropical decomposition to obtain the below observation.
@@
Let $f$ define a Newton Plane Curve singularity such that $(1,1)\notin supp(f)$. Then for every $\alpha\in T(f)$, $\nu_{\alpha}(f)>|\alpha|$ unless $\alpha\in\mathbf{R}_{\geq 0}\cdot(1,1)$ with $f$ having two branches with $T(f)=\mathbf{R}_{\geq 0}\cdot(1,1)$.
@@
Since we can always consider the symmetry $(x,y)\mapsto(y,x)$, let $\alpha=(m,n)$ with $m\geq n\geq 1$ and $gcd(m,n)=1$. Suppose
@@
where $f_{\alpha}$ denotes the product of branches having tropical variety $\mathbf{R}_{\geq 0}\cdot\alpha$. Let’s show $f=f_{\alpha}$. Otherwise, $t\geq 2$ where $t$ is the number of semi-lines in the tropical variety $T(f)$. Denote by $f=f_{1}...f_{t}$ an oriented tropical decomposition of $f$. We have $f_{\alpha}\in\{f_{1},...,f_{t}\}$. There exists $S=(a,b)$ a vertex in the Newton polygon of $f$ with $a,b\geq 1$ (because $t>1$) and $\nu_{\alpha}(f)=\alpha\cdot S$. Set $\eta=(1,1)-S$
$0\leq|\alpha|-\nu_{\alpha}(f)=\alpha\cdot\eta$
Note that $\alpha_{1},\alpha_{2}\geq 1$. Since $a,b\geq 1$ then $\alpha\cdot\eta=0$ but $\alpha$ is in the first positive quadrant $\mathbf{R}^{2}_{>0}$ and $\eta\in\mathbf{R}^{2}_{\leq 0}$ so they can’t be orthogonal unless $\eta=0$ ie $S=(1,1)\in supp(f)$, a contradiction. So $t=1$ and $f=f_{\alpha}$.
@@
where $r$ is the number of branches of $f$. According to Observation 1, $n=1$ then
$1\leq r\leq 1+\frac{1}{m}$
If $m>1$ then $r=1$ we can write in Weierstrass form
@@
Thus $f$ is smooth, a contradiction with $f$ defining a singularity. So $r=2$ and $m=1$. So, we can write in Weierstrass form after doing a change of variables
f=(yx+)(ycx+)f=(y-x+...)(y-cx+...)
Thus, $|\alpha|=\nu_{\alpha}(f)=2$ with $\alpha=(1,1)$. This concludes the proof of our observation. ∎
In the below lemma, we write an expression for $H$ (which was introduced at the beginning of the section). There are many cases to consider:
@@
Let $f\in K[x,y]$ be a Newton non-degenerate polynomial whose oriented tropical decomposition is denoted $f=f_{1}...f_{t}$. Then $H$ is the summation
@@
where $m\geq 1$, $Cont^{m}(\mathcal{C})^{0}\neq\varnothing$ and $\beta\in B^{*}_{m}$. Also, $H\in\mathbf{Z}(u,v)$. Moreover, if $S_{1},...,S_{t+1}$ are the vertices of the Newton polygon of $f$ from left to right, we have
  • If $t=1$ and $T(f)=\mathbf{R}_{\geq 0}(1,1)$ then $H=\frac{u^{2}v^{s}}{1-u^{2}v^{s}}$ where the Newton polygon of $f$ is the segment joining $(s,0)$ and $(0,s)$,
  • If $t=1$ and $T(f)>\mathbf{R}_{\geq 0}(1,1)$ then $H=H_{T}+H_{\overset{\circ}{\sigma_{1}}}+H_{\overset{\circ}{\sigma_{2}}}$ where
    $H_{T}=\frac{u^{|\alpha|}v^{\alpha\cdot S}}{1-u^{|\alpha|}v^{\alpha\cdot S}}$
    where $T(f)=\mathbf{R}_{\geq 0}\cdot\alpha$, $gcd(\alpha_{1},\alpha_{2})=1$ and $S=(s,0)$ is a vertex of the Newton polygon of $f$,
    $H_{\overset{\circ}{\sigma_{1}}}=\frac{1}{1-u^{|\alpha|}v^{\alpha\cdot S}}\sum_{i=1}^{p}\sum_{j=D_{i-1}+1}^{D_{i}}u^{i+j}v^{js}$
@@
    where $SC(\alpha_{2}/\alpha_{1})=[[d_{1},...,d_{p}]]$ and $D_{i}=\sum_{j=1}^{i}d_{j}$, $D_{0}=0$.
  • If $t>1$, set $T_{i}=T(f_{i})=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$ with $T_{1}<...<T_{t}$ and $T_{1}=\mathbf{R}_{\geq 0}(1,1)$ then
@@
    where $H_{T_{i}}=\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}$ and for $2\leq i\leq t$, we have:
@@
    where $SC(\alpha_{2,t}/\alpha_{1,t})=[[d_{1t},...,d_{p_{t},t}]]$, $D_{i,t}=\sum_{j=1}^{i}d_{j,t}$, $D_{0,t}=0$ and $S_{t+1}=(s_{t+1},0)$.
  • If $t>1$, set $T_{i}=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$ with $\mathbf{R}_{\geq 0}(1,1)<T_{1}<...<T_{t}$ then
@@
    with $H_{T_{i}}=\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}$, and for $2\leq i\leq t+1$, $H_{\overset{\circ}{\sigma_{i}}}$ is as the previous point and
@@
    $H_{T_{i}}$ are as before, $H_{\overset{\circ}{\sigma_{i}}}$ as before for $2\leq i\leq t+1$ and finally,
@@
This comes from the fact that $H$ has a decomposition:
@@
where $H_{T_{i}}$ is the sum over the lattice points who lie on $T_{i}\cap\mathbf{N}^{2}$ and $H_{\overset{\circ}{\sigma_{i}}}$ is the sum over the lattice points who lie in the interior of the cone $\sigma_{i}=T_{i-1}+T_{i}$. We have always:
@@
since $\alpha_{i-1}\cdot S_{i}=\alpha_{i-1}\cdot S_{i-1}$ for $\alpha_{i}\in\sigma_{i}$. For $H_{\overset{\circ}{\sigma_{1}}}$ and $H_{\overset{\circ}{\sigma_{t+1}}}$, there are multiple cases depending on whether $T(f_{1})\geq\mathbf{R}_{\geq 0}(1,1)$ or $T(f_{1})<\mathbf{R}_{\geq 0}(1,1)<T(f_{t})$.
  • if $T(f_{1})=\mathbf{R}_{\geq 0}(1,1)$ then $H_{\overset{\circ}{\sigma}_{1}}=0$ ;
  • if $T(f_{1})>\mathbf{R}_{\geq 0}(1,1)$ and $t=1$ then choose $\alpha=(p,q)$ the generator of the tropical variety of $f$ such that $gcd(p,q)=1$ $SC(q/p)=[[d_{1},...,d_{p}]]$. Also,
@@
    where $S_{1}=(0,s_{1})$ is the first vertex of the Newton polygon of $f$. As for $H_{\overset{\circ}{\sigma}_{2}}$ we have:
@@
    where $V=\{(1,d_{1}+1),(2,d_{1}+d_{2}+1),...,(p-1,d_{1}+...+d_{p-1}+1)\}$.
  • if $T(f_{1})>\mathbf{R}_{\geq 0}(1,1)$ and $t>1$ then $H_{\overset{\circ}{\sigma}_{1}}$ (resp. $H_{\overset{\circ}{\sigma}_{t+1}}$) is computed as $H_{\overset{\circ}{\sigma}_{1}}$ (resp. $H_{\overset{\circ}{\sigma}_{2}}$) in the previous point but with replacing $T(f)$ by $T(f_{1})$ (resp. by $T(f_{t})$) and $SC(q/p)$ by $SC(\alpha_{1,2}/\alpha_{1,1})$ (resp. $SC(\alpha_{t,2}/\alpha_{t,1})$).
  • if $T(f_{1})<\mathbf{R}_{\geq 0}(1,1)<T(f_{t})$, then $H_{\overset{\circ}{\sigma}_{t+1}}$ is computed as $H_{\overset{\circ}{\sigma}_{t+1}}$ in the previous point. As for $H_{\overset{\circ}{\sigma}_{1}}$ it is computed as such:
@@
    where $U^{\prime}$ is the same set as $U$ but where axis and ordinates of each lattice point are flipped.
When we compute each of $H_{\overset{\circ}{\sigma}_{1}}$ and $H_{\overset{\circ}{\sigma}_{t+1}}$ in detail we find the expression written in the statement of the theorem. The case where $T(f_{1})>T(f_{t})$ doesn’t exist because we have supposed

### v3

@@ 8. Generating Series
In this section, we introduce the generating series $G:=G(f)$ associated to a Newton non-degenerate power series $f\in K[[x,y]]$. If $\mathcal{C}=\{f=0\}$ denotes its associated plane curve singularity, then we define
@@
where $m\geq 1$ and $K$ is an irreducible component in $\text{Cont}^{m}(\mathcal{C})^{0}\neq\varnothing$. Now given a sum $S$, then a subsummation of $S$ is a sum whose terms form a subset of the terms of $S$.
@@
  • $H$ is the subsummation of $G$ with $K$ being of the form $\mathcal{H}^{\beta}_{\infty}$, and
  • $R$ is the rest of the sum
It is worthy to note that here we have used the letter $H$ because the $K$s appearing in $H$ are all irreducible components having hyperplane coordinates and R=GHR=G-H is just the rest of the sum.
Let us start with the following two different examples.
@@
Let $f=(y^{2}-x^{3})(y^{2}-2x^{3})$ whose tropical variety $T(f)=\mathbf{R}_{\geq 0}(2,3)$ is clearly a semi-line. Set $G=G(f)$ and let
@@
where $m\geq\nu_{\alpha}(f)$ and $\alpha\in T(f)$. Since $\alpha=s\cdot(2,3)$ for $s\in\mathbf{N}$, then
$R=2u\sum_{s=1}^{\infty}\sum_{m=12s}^{\infty}u^{5s+m-12s}v^{m}=\frac{2u}{1-uv}\sum_{s=1}^{\infty}u^{5s}v^{12s}=\frac{2u}{1-uv}\frac{u^{5}v^{12}}{1-u^{5}v^{12}}.$
Now, set $S_{1}=(0,4)$, $S_{2}=(6,0)$ and let $S^{*}_{i}$ be the corresponding cone to $S_{i}$ in the dual Newton fan. In addition, let
@@
where $\beta_{i}\in B^{*}_{m}\cap\overset{\circ}{S^{*}_{i}},m\geq 1$. In order to describe where $\beta_{i}$ lies, note that
@@
for $\lambda\in T(f)\cap\mathbf{N}_{0}^{2}$, and so we get $\beta_{2}-\lambda\in\{(1,2)\}$ for the same $\lambda=(2s,3s)$. This means that
@@
$H=\frac{u^{5}v^{12}}{1-u^{5}v^{12}}+\frac{u^{2}v^{4}}{1-u^{5}v^{12}}+\frac{u^{4}v^{8}}{1-u^{5}v^{12}}+\frac{u^{3}v^{6}}{1-u^{5}v^{12}}\in\mathbf{Z}(u,v).$
Thus, $G=H+R\in\mathbf{Z}(u,v)$.
@@
Consider $f=f_{1}f_{2}=(y^{2}-x^{3})(y^{3}-x^{5})$ whose tropical variety is
$T(f)=\mathbf{R}_{\geq 0}(2,3)\cup\mathbf{R}_{\geq 0}(3,5).$
Now we compute
@@
where $\alpha_{i}\in T(f_{i})$, $m_{i}\geq\nu_{\alpha_{i}}(f)$. In this case,
$R_{1}=u\sum_{s=1}^{\infty}\sum_{m=0}^{\infty}u^{5s+m}v^{m+15s}=\frac{u}{1-uv}\frac{u^{5}v^{15}}{1-u^{5}v^{15}},$
and
$R_{2}=\frac{u}{1-uv}\frac{u^{8}v^{24}}{1-u^{8}v^{24}}.$
Next, let $S_{1}=(0,5)$, $S_{2}=(3,3)$ and $S_{3}=(8,0)$ be the vertices of the Newton polygon from left to right. As earlier, let $S_{i}^{*}$ be the corresponding cone to $S_{i}$ in the dual Newton fan. Then we can wrote
@@
where $\beta\in B^{*}_{m},m\geq 1$ and $\sigma_{2}=\mathbf{R}_{\geq 0}(2,3)+\mathbf{R}_{\geq 0}(3,5)=S^{*}_{2}$. By computing the geometric series, we obtain
$D:=H-\tilde{H}=\frac{u^{5}v^{15}}{1-u^{5}v^{15}}\frac{u^{8}v^{24}}{1-u^{8}v^{24}}\in\mathbf{Z}(u,v).$
Next, we compute
@@
where $H^{*}$ is the subsummation of $H$ whose $\beta$ lies in the first and the last cone determined by the orientation of the tropical decomposition. Then, $\tilde{H}-H^{*}=\frac{u^{5}v^{15}}{1-u^{5}v^{15}}+\frac{u^{8}v^{24}}{1-u^{8}v^{24}}\in\mathbf{Z}(u,v)$ so that
@@
where $\beta_{i}\in\overset{\circ}{S_{i}^{*}}\cap B^{*}_{m},m\geq 1$, $S_{1}=(0,5)$ and $S_{3}=(8,0)$. Therefore,
@@
which leads to
$H^{*}=\frac{u^{2}v^{5}+u^{4}v^{10}}{1-u^{5}v^{15}}+\frac{u^{3}v^{8}+u^{6}v^{16}}{1-u^{8}v^{24}}\in\mathbf{Z}(u,v).$
Thus, $H\in\mathbf{Z}(u,v)$ and $G=H+R$ as well.
In the next lemma, we shall write an expression for $R$ (which was introduced at the beginning of the section).
@@
Let $f\in K[x,y]$ be Newton non-degenerate whose tropical decomposition is written as $f=f_{1}...f_{t}$. Also, for $1\leq i\leq t$, let $f_{i}=\prod_{j=1}^{r_{i}}f_{i,j}$. Then,
@@
where $1\leq i\leq t$, $1\leq j\leq r_{i}$, $\alpha\in T(f_{i})$ and $m\geq\nu_{\alpha}(f)$. Moreover, $R\in\mathbf{Z}(u,v)$ and more explicitly,
@@
Set $R_{i}=\sum_{j,\alpha_{i},m}u^{1+|\alpha_{i}|+m-\nu_{\alpha_{i}}(f)}v^{m}$ in such a way that $R=\sum_{i}R_{i}$. Since $\nu_{\alpha}(f)=\alpha\cdot S_{i}$ (because $\alpha\in T(f_{i})$), then we obtain:
@@
Now let $f=f_{1}...f_{t}$ be a tropical decomposition of a Newton non-degenerate polynomial. Then, the decomposition $f=f_{1}...f_{t}$ is oriented in such a way that for all $i<j$,
@@
i.e. $T(f_{j})$ is ”above” $T(f_{i})$ for which we write $T(f_{j})>T(f_{i})$. Making use of this new concept of oriented tropical decomposition, we obtain the following observation.
@@
Let $f$ define a Newton Plane Curve singularity such that $(1,1)\notin supp(f)$. Then, for every $\alpha\in T(f)$, we have $\nu_{\alpha}(f)>|\alpha|$ unless $\alpha\in\mathbf{R}_{\geq 0}(1,1)$ with $f$ having two branches and $T(f)=\mathbf{R}_{\geq 0}\cdot(1,1)$.
@@
Due to the symmetry $(x,y)\mapsto(y,x)$, then without loss of generality we assume that $\alpha=(m,n)$ with $m\geq n\geq 1$ and $\gcd(m,n)=1$. Suppose that
@@
where $f_{\alpha}$ denotes the product of branches having tropical variety $\mathbf{R}_{\geq 0}\cdot\alpha$. Now we show that $f=f_{\alpha}$. If not, then $t\geq 2$ where $t$ is the number of semi-lines in the tropical variety $T(f)$. Denote by $f=f_{1}...f_{t}$ an oriented tropical decomposition of $f$. Since $f_{\alpha}\in\{f_{1},...,f_{t}\}$, then there exists $S=(a,b)$ a vertex in the Newton polygon of $f$ with $a,b\geq 1$ (as $t>1$) and $\nu_{\alpha}(f)=\alpha\cdot S$. Next, set $\eta=(1,1)-S$ then
$0\leq|\alpha|-\nu_{\alpha}(f)=\alpha\cdot\eta.$
Note that as $\alpha_{1},\alpha_{2}\geq 1$, and $a,b\geq 1$, then this last inequality implies that $\alpha\cdot\eta=0$. However, $\alpha$ being in the first positive quadrant $\mathbf{R}^{2}_{>0}$ and $\eta$ being in the third quadrant $\mathbf{R}^{2}_{\leq 0}$, they can not be orthogonal unless $\eta=0$ and then $S=(1,1)\in supp(f)$ which is a contradiction. Therefore, $t=1$ and $f=f_{\alpha}$. Observe that
@@
where $r$ is the number of branches of $f$. In view of Observation 1, $n=1$ which in turn implies that
$1\leq r\leq 1+\frac{1}{m}.$
If $m>1$, then $r=1$ and we can write in Weierstrass form
@@
and this leads to $f$ being smooth, which contradicts the fact that $f$ defines a singularity. Therefore, $r=2$ and $m=1$. So that after performing a change of variables, we can write in Weiersrass form
f=(yx+)(ycx+).f=(y-x+...)(y-cx+...).
Thus, $|\alpha|=\nu_{\alpha}(f)=2$ with $\alpha=(1,1)$, and this concludes the proof of our observation. ∎
In the next lemma, we shall give an expression for $H$ (which was introduced at the beginning of the section). It is worth recalling here the following:
@@
Let $f\in K[x,y]$ be a Newton non-degenerate polynomial whose oriented tropical decomposition is denoted $f=f_{1}...f_{t}$. Then $H$ is given by
@@
where $m\geq 1$, $Cont^{m}(\mathcal{C})^{0}\neq\varnothing$ and $\beta\in B^{*}_{m}$. Also, $H\in\mathbf{Z}(u,v)$. Moreover, if $S_{1},...,S_{t+1}$ are the vertices of the Newton polygon of $f$ from left to right, then, the following holds.
  • If $t=1$ and $T(f)=\mathbf{R}_{\geq 0}(1,1)$, then $H=\frac{u^{2}v^{s}}{1-u^{2}v^{s}}$ where the Newton polygon of $f$ is the line=segment joining $(s,0)$ and $(0,s)$,
  • If $t=1$ and $T(f)>\mathbf{R}_{\geq 0}(1,1)$, then $H=H_{T}+H_{\overset{\circ}{\sigma_{1}}}+H_{\overset{\circ}{\sigma_{2}}}$ with
    $H_{T}=\frac{u^{|\alpha|}v^{\alpha\cdot S}}{1-u^{|\alpha|}v^{\alpha\cdot S}},$
    and where $T(f)=\mathbf{R}_{\geq 0}\cdot\alpha$, $\gcd(\alpha_{1},\alpha_{2})=1$ and $S=(s,0)$ is a vertex of the Newton polygon of $f$ as well as
    $H_{\overset{\circ}{\sigma_{1}}}=\frac{1}{1-u^{|\alpha|}v^{\alpha\cdot S}}\sum_{i=1}^{p}\sum_{j=D_{i-1}+1}^{D_{i}}u^{i+j}v^{js},$
@@
    where $D_{0}=0$, $D_{i}=\sum_{j=1}^{i}d_{j}$ and the $d_{i}$’s come from $SC(\alpha_{2}/\alpha_{1})=[[d_{1},...,d_{p}]]$.
  • If $t>1$, then by setting $T_{i}=T(f_{i})=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$ with $T_{1}<...<T_{t}$ and $T_{1}=\mathbf{R}_{\geq 0}(1,1)$ we obtain
@@
    where $H_{T_{i}}=\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}$ and
@@
    where $S_{t+1}=(s_{t+1},0)$, $D_{0,t}=0$, $D_{i,t}=\sum_{j=1}^{i}d_{j,t}$ and the $d_{i}$’s again come from $SC(\alpha_{2,t}/\alpha_{1,t})=[[d_{1t},...,d_{p_{t},t}]]$.
  • If $t>1$, then by setting $T_{i}=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$ with $\mathbf{R}_{\geq 0}.(1,1)<T_{1}<...<T_{t}$ we get
@@
    with $H_{T_{i}}=\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}$, and $H_{\overset{\circ}{\sigma_{i}}}$ is given by the previous case for all $2\leq i\leq t+1$ as well as
@@
    where $H_{T_{i}}$, $H_{\overset{\circ}{\sigma_{i}}}$ are as before for all $2\leq i\leq t+1$ and with
@@
The first part comes from the fact that $H$ has the decomposition:
@@
where $H_{T_{i}}$ is the sum over the lattice points which lie on $T_{i}\cap\mathbf{N}^{2}$ and $H_{\overset{\circ}{\sigma_{i}}}$ is the sum over the lattice points which lie in the interior of the cone $\sigma_{i}=T_{i-1}+T_{i}$. More explicitly,
@@
as $\alpha_{i-1}\cdot S_{i}=\alpha_{i-1}\cdot S_{i-1}$ for $\alpha_{i}\in\sigma_{i}$. For $H_{\overset{\circ}{\sigma_{1}}}$ and $H_{\overset{\circ}{\sigma_{t+1}}}$, there are multiple cases depending on whether $T(f_{1})\geq\mathbf{R}_{\geq 0}(1,1)$ or $T(f_{1})<\mathbf{R}_{\geq 0}(1,1)<T(f_{t})$.
  • If $T(f_{1})=\mathbf{R}_{\geq 0}(1,1)$, then $H_{\overset{\circ}{\sigma}_{1}}=0$.
  • If on the other hand $T(f_{1})>\mathbf{R}_{\geq 0}(1,1)$ and $t=1$, then choose $\alpha=(p,q)$ as the generator of the tropical variety of $f$ such that $\gcd(p,q)=1$ and $SC(q/p)=[[d_{1},...,d_{p}]]$. In addition,
@@
    where $S_{1}=(0,s_{1})$ is the first vertex of the Newton polygon of $f$. As for $H_{\overset{\circ}{\sigma}_{2}}$, the following holds
@@
    with $V=\{(1,d_{1}+1),(2,d_{1}+d_{2}+1),...,(p-1,d_{1}+...+d_{p-1}+1)\}$.
  • If $T(f_{1})>\mathbf{R}_{\geq 0}(1,1)$ and $t>1$, then $H_{\overset{\circ}{\sigma}_{1}}$ (resp. $H_{\overset{\circ}{\sigma}_{t+1}}$) can be computed in the same way as $H_{\overset{\circ}{\sigma}_{1}}$ (resp. $H_{\overset{\circ}{\sigma}_{2}}$) in the previous case but with replacing $T(f)$ by $T(f_{1})$ (resp. by $T(f_{t})$) and $SC(q/p)$ by $SC(\alpha_{1,2}/\alpha_{1,1})$ (resp. $SC(\alpha_{t,2}/\alpha_{t,1})$).
  • If $T(f_{1})<\mathbf{R}_{\geq 0}(1,1)<T(f_{t})$, then $H_{\overset{\circ}{\sigma}_{t+1}}$ can be found in the same way as $H_{\overset{\circ}{\sigma}_{t+1}}$ in the previous case too. As for $H_{\overset{\circ}{\sigma}_{1}}$ it is given by:
@@
    where $U^{\prime}$ is the same set as $U$ but where axes and ordinates of each lattice point are flipped.
Now a simple check shows that each of $H_{\overset{\circ}{\sigma}_{1}}$ and $H_{\overset{\circ}{\sigma}_{t+1}}$ are given as in the statement of the theorem. Finally, observe that the case where $T(f_{1})>T(f_{t})$ can not occur since we have assumed that